OpenStudy (anonymous):

how to do I change lny = -1/2lnx + 6.8 into the form y = Ax^b ???

OpenStudy (amistre64):

stick an e on it

OpenStudy (amistre64):

some use of ( ) to clean up the equation might be useful ...

OpenStudy (anonymous):

ummm...... can you expand on that???? lets say I have this: lny = lnx^-1/2 + 6.8 Now what??

OpenStudy (anonymous):

lny=lnx^(-0.5)+6.8 y=x^-0.5*e^6.8 Let e^6.8 be A y=Ax^-0.5

OpenStudy (amistre64):

bases and logs are inverses, they cancel each other out

OpenStudy (amistre64):

lny = lnx^-1/2 + 6.8 e^(lny = lnx^-1/2 + 6.8 ) y = x^-1/2 * e^(6.8)

OpenStudy (anonymous):

ohhhhhhh, zed, that makes sense.

OpenStudy (anonymous):

and you too, amistre64 :D

OpenStudy (amistre64):

\begin{array}l\color{red}{\text{y}}\color{orange}{\text{a}}\color{#9c9a2e}{\text{y}}\color{green}{\text{!}}\color{blue}{\text{!}}\color{purple}{\text{ }}\color{purple}{\text{:}}\color{red}{\text{)}}\color{orange}{\text{}}\end{array}

OpenStudy (anonymous):

^lol